# Galaxy Groups in the presence of Cosmological Constant: Re-Mapping the close-by galaxies

### v1

@@ Abstract
Boundaries of galaxy groups and clusters are defined by the interplay between the Newtonian attractive force and the local repulsion force caused by the expansion of the Universe. This research extends the definition of zero radial acceleration surface (ZRAS) and the turnaround surface (TS) for a general distribution of the masses in an expanding background, governed by the cosmological constant. We apply these definitions for different galaxy groups in the local Universe, mapping these groups up to ten megaparsec. We discuss the dipole and the quadrupole rate for the Local Group of Galaxies and the implementations for Hubble diagram correction and galaxy groups viralization. With these definitions, we indicate the surfaces showing the interplay between the local expansion vs the local Newtonian attraction for galaxy groups in the local Universe. The results show that it is important to include the cosmological constant in analyzing the Cosmic Flow of the local universe.
@@ I Introduction
Dark energy is an unknown type of energy that influences the Universe on its largest scales Peebles and Ratra (2003). Its primary impact is to propel the accelerating expansion of the Universe. The first observational evidence for dark energy emerged from supernova measurements Scolnic et al. (2018) and later was confirmed by the Cosmic Microwave Background radiation measurements Aghanim et al. (2020) and the Baryon Acoustic Oscillations. Beyond its impact on large scales, dark energy exerts a significant influence on the Local Universe Courtois et al. (2013). Galaxy groups can be characterized by the radius of decoupling from cosmic expansion Tully (2015). In an isolated system, the dark energy creates a local repulsion force resisting the Newtonian attraction:
@@
where $r$ is the separation, $M$ is the total mass of the binary system, $G$ is the Newtonian gravitational constant, $c$ is the speed of light and $\Lambda$ is the cosmological constant.
It is possible to infer the cosmological constant from the Cosmological Observable via:
$\Lambda=3\Omega_{\Lambda}\left(H_{0}/c\right)^{2}\;,$ (2)
where $H_{0}$ the Hubble parameter that estimates the current expansion rate, and $\Omega_{\Lambda}$ is the dark energy density rate. From the value of the Hubble constant derived by the Planck collaboration, the $\Lambda$ is $\Lambda_{\text{CMB}}=\left(1.08\pm 0.02\right)\cdot 10^{-52}\,m^{-2},$ Aghanim et al. (2020), while the value from the SH0ES data is: $\Lambda_{\text{SH0ES}}=\left(1.26\pm 0.05\right)\cdot 10^{-52}\,m^{-2}$ Brout et al. (2022) with a $3\sigma$ difference.
Since the dark energy contribution depends on the mass density, we define two useful surfaces: the Zero Radial Acceleration Surface (ZRAS) and the Turnaround Surface (TS). These two definitions are compatible for the spherical case, and in this paper we generalize them for a general density structure.
@@
    $\bar{r}_{0}\equiv\sqrt[3]{\frac{3GM}{\Lambda c^{2}}},$ (3)
    which is called the zero gravity radius. In astronomical units:
    $\bar{r}_{0}\approx 1.11\,\text{Mpc}\,\left(\frac{M}{10^{12}\,M_{\odot}}\right)^{1/3}\left(\frac{67\,km/s/Mpc}{H_{0}}\right)^{2/3},$ (4)
    where $M_{\odot}$ is the solar mass. There are different methods to measure the zero gravity radius through the local Hubble expansion Del Popolo and Chan (2022) or by estimating the points at which the sphere breaks away from the general expansion and reaches a maximum radius Pavlidou and Tomaras (2014); Faraoni et al. (2015); Faraoni (2016); Giusti and Faraoni (2021). For the spherical case, the zero gravity radius is sufficient, but for a general density distribution one has to use the ZRAS, which reads Dolgachev et al. (2003):
@@
  • Turnaround Surface (TS) – In statistical mechanics, the virial theorem provides a general equation that relates the average over time of the total kinetic energy of a stable, closed system of discrete particles, bound by a conservative force Lahav et al. (1991); Mota and van de Bruck (2004); Wang (2006); Chernin et al. (2012). The conventional virial theorem can be modified by an extra term due to the contribution of the dark energy, besides the Newtonian attraction. By imposing that all the kinetic energy of the system is positive ($K>0$) one gets the condition Nowakowski et al. (2002):
    $\frac{U_{G}}{2}<U_{\Lambda}.$ (6)
@@
    The condition for positive kinetic energy is extended for a general density case and gives the definition for the viralization surface. In the standard case where the dispersion of the velocities is known, the virial theorem gives the lowest possible radii for a closed system. In this case it is possible to calculate the value of the average spherical radii. Correspondingly, there exists a surface that obeys the virial theorem for the particles inside the surface.
Galaxy groups can be characterized by the radius of decoupling from cosmic expansion Tully (2015), however this definition was made without the presence of dark energy. In this paper, we extend these definitions from a spherical symmetric case to a general density via moltipole expansion and we define the corresponding ZRAS and the TS. The structure of the paper is as follows: Section II reviews the theoretical framework for a spherical density model with dark energy. Section III extends the analysis for a galaxy pair. Section IV shows the applications for galaxies in the local Universe and maps the local galaxy complexes correspondingly. Section VI summarizes the results of this paper.
@@ II Point Particle and Spherical Symmetry
Zero Acceleration Surface.
@@
Turnaround Surface.
The virial theorem provides a general equation that relates the average over time of the total kinetic energy of a stable system of discrete particles, bound by a conservative force, to that of the total potential energy of the system. The virial theorem relates the average a potential energy with a power law form $U\sim r^{n}$, via the relation Hill (1961):
$\langle K\rangle=\frac{n}{2}\langle U\rangle,$ (9)
where $K$ is the kinetic energy of the system, $U$ is the potential energy of the system and the averages are taken over long time. The average yields:
$K+\frac{1}{2}U_{G}-U_{\Lambda}=0,$ (10)
@@
$K=\frac{1}{2}\sum m_{i}v_{i}^{2},$
$U_{G}=G\sum_{i<j}\frac{m_{i}m_{j}}{r_{ij}},$
$U_{\Lambda}=\frac{1}{6}\Lambda c^{2}\sum_{i}m_{i}r_{i}^{2}.$ (11)
@@
As a simple model for the galaxy groups in the local Universe, it is assumed that most of the mass is concentrated in the massive galaxies, and the other dwarf galaxies are test particles with mass $m$. For a test particles obeying the Eqs. (11) with the condition (6), one finds
$r<\bar{r}_{0},$ (12)
where $\bar{r}_{0}$ is the ZRAS for the massive galaxy $M$. It means that for a closed system within the zero radial acceleration surface. The estimate indicates that gravitation is stronger than repulsion within the volume of a group, making the existence of gravitationally bound systems with finite orbits possible. Based on this condition, it was suggested in Membrado and Pacheco (2016) to use surface term effects as mass estimators. From this inequality, it is clear that this gives an upper value for the mass inside the spherical surface.
III Galaxy Pair and Axial Symmetry
@@ III.1 The surfaces
where $m_{1,2}$ are the fractional masses from the total mass:
$m_{1,2}=M_{1,2}/(M_{1}+M_{2}).$ (15)
Fig. (1) shows an illustration of this system. A test particle around these galaxies feels the potential
@@
Figure 1: Stream lines for a galaxy pair with masses $M_{1,2}$ located at $\left(0,z_{1,2}\right)$. The stream lines shows the attractor for the gravitational force inside the surface described by the Eq. (23), vs the repulsion field outside
Critical points.
@@
$F_{z}=\frac{\Lambda c^{2}z}{3}-\frac{GM_{1}(z-z_{1})}{\left(\rho^{2}+(z-z_{1})^{2}\right)^{3/2}}-\frac{GM_{2}(z-z_{2})}{\left(\rho^{2}+(z-z_{2})^{2}\right)^{3/2}},$ (17a)
$F_{\rho}=\frac{\Lambda c^{2}\rho}{3}-\frac{GM_{1}\rho}{\left(\rho^{2}+(z-z_{1})^{2}\right)^{3/2}}-\frac{GM_{2}\rho}{\left(\rho^{2}+(z-z_{2})^{2}\right)^{3/2}}.$ (17b)
@@
  • Points A, C $(0,\pm z_{0})$:
    The force at these points is zero, ${\bf F}_{A,B}={0}$. For $\rho=0$, we get $F_{\rho}=0$ from Eq. (17b). Imposing $F_{z}=0$ in Eq. (17a) gives:
@@
    By normalizing the Eq. (18) using the zero radial acceleration for the total mass $\bar{r}_{0}=\bar{r}_{0}\left(M\right)$, the equation is simplified to:
    $\frac{m_{1}}{(1-m_{2}z_{12}/z_{0})^{2}}+\frac{m_{2}}{(1+m_{1}z_{12}/z_{0})^{2}}=\left(\frac{z_{0}}{\bar{r}_{0}}\right)^{3}.$ (19)
@@
    $\frac{GM_{1}}{\left(\rho_{0}^{2}+z_{1}^{2}\right)^{3/2}}+\frac{GM_{2}}{\left(\rho_{0}^{2}+z_{2}^{2}\right)^{3/2}}=\frac{\Lambda c^{2}}{3},$ (20)
    and with the normalization of the radial acceleration surface is:
    $\frac{m_{1}}{\left(1+m_{2}^{2}z_{12}^{2}/\rho_{0}^{2}\right)^{3/2}}+\frac{m_{2}}{\left(1+m_{1}^{2}z_{12}^{2}/\rho_{0}^{2}\right)^{3/2}}=\left(\frac{\rho_{0}}{\bar{r}_{0}}\right)^{3}.$ (21)
@@
Radial acceleration surface.
@@
$\frac{\Lambda c^{2}r^{2}}{3}=\frac{GM_{1}\left(r^{2}-zz_{1}\right)}{\left(r^{2}+z_{1}(z_{1}-2z)\right)^{3/2}}+\frac{GM_{2}\left(r^{2}-zz_{2}\right)}{\left(r^{2}+z_{2}(z_{2}-2z)\right)^{3/2}}.$ (23)
The equation can be rewritten as:
$\left(\frac{r}{\bar{r}_{0}}\right)^{3}=\frac{m_{1}\left(1-zz_{1}/r^{2}\right)}{\left(1-2zz_{1}/r^{2}+z_{1}/r^{2}\right)^{3/2}}+\mathcal{O}\left(1\rightarrow 2\right).$ (24)
The terms $zz_{i}/r^{2}$ and $z_{i}/r^{2}$ are dimensionless. Fig. (1) shows in dashed black line the solution for this curve. The plot shows that this curve differentiates the gravitational attraction from the global cosmic expansion.
Turnaround surfaces.
@@ III.2 Expansion
To study the properties of the potential, one can expand the potential around $z_{i}/r\ll 1$, assuming that the observer is distant from the group. The approximation gives:
@@
$\sigma_{\text{Quad}}\equiv\frac{m_{1}}{2}z_{1}^{2}+\frac{m_{2}}{2}z_{2}^{2}.$ (29)
Since we work in CoM coordinates, i.e. CoM is at the origin (Eq. 13), the dipole of the system is zero Peñarrubia et al. (2016). The dipole, octuple, and all multipoles with odd powers in $r^{i}$ defined in the center of mass of any system should be equal to zero. In order to quantify the contribution of the quadropole, we can test the dimensionless quantity
@@
IV Mapping the Local Universe
Figure 2: Gradient lines (in purple) for the Local Group of galaxies including the Milky Way (with $M_{1}\approx 10^{12}\,M_{\odot}$) and M31 ($M_{2}\approx 2\cdot 10^{12}\,M_{\odot}$) with a distance of $0.77\,Mpc$, located at $(0,0.77\cdot 2/3)$ and $(0,-0.77/3)$, respectively. The center of mass is at the origin of the system of coordinates. The streamlines show the areas for the domination of the Newtonian attraction vs the cosmic repulsion. The intersection of the ZRAS and the TS is presented by the dashed black lines and the dashed red lines respectively. The dwarf galaxies around the Milky Way (blue) and M31 (purple) are also added as points. The TS is independent of the velocity distribution and compatible with most dwarf galaxies, in addition to a few outliers.
@@ IV Mapping the Local Universe
Milky-Way Mass $[0.5,1.3]\cdot 10^{12}\,M_{\odot}$ Wang et al. (2020)
Andromeda Mass $[1.0,2.0]\cdot 10^{12}\,M_{\odot}$ Sawala et al. (2023)
@@
Table (2) shows the data we consider for the LG of galaxies to estimate the characteristic surfaces. Table (2) shows the different values for the zero acceleration surface (monopole), the critical points ZRAS and TS. We can see from the figure that the surfaces is non-spherical, where one edge is bigger then $\bar{r}_{0}$ and the other edge is smaller then $\bar{r}_{0}$. The intersection of this surface with the symmetry plane (as in Fig. (2)) is represented by a slightly asymmetric oval that lies between two circles, where $z_{1}/\bar{r}_{0}\approx 33\%$ and $z_{2}/\bar{r}_{0}\approx 20\%,$ are the partial corrections for the galaxy locations over the monopole ZRAS, with an error less then one percent. For the LG, the corresponding quadrupole correction is $\tilde{\sigma}_{\text{Quad}}^{(\text{LG})}<3.2\%$. Therefore, treating the quadrupole as a perturbation over the monopole and Dark Energy parts is a valid assumption.
Figure 3: Two projections of the Local Group (up to 10 Mpc) and their immediate vicinity. The dashed curves show the gravitational bond (with dark energy) for massive galaxies. The figure includes the Local Group, the Cen-A/M83 complex, the M81-M82 group, and the NGC 253 Group. The colors of the dwarf galaxies around show the velocities with respect to the CMB frame as taken from Cosmic Flow Hoffman et al. (2023).
@@
One Galaxy – The simplest case is one dominant giant galaxy and some dwarf galaxies around. The NGC 253 system and the IC-342 group and M101 galaxies seem to satisfies this condition. In this case, the spherical symmetry is a good approximation, unless there is a distribution of dark matter around the system that breaks this symmetry. Therefore, around these giant galaxies the TS looks as a perfect circles, despite that the dark matter distribution could give additional quadrupole corrections.
Galaxy Pair – The LG has two main components, the MW and M31. In fact, this binary structure has been suggested as a possible source for the ejection of dwarf galaxies. However, the zero-gravity surface calculated for the LG is almost spherically symmetric at the relevant distances. The CenA & M83 Complex system seems to be dominated by two giant galaxies: Centaurus A (NGC 5128) and M83 (NGC 5236), where the other galaxies are dwarf galaxies around these. Another pair is the M81-M82 Complex, where the M81 group is a galaxy group in the constellations Ursa Major and Camelopardalis. For these systems, the TS gives an egg shape as discussed in the LG, but since the quadropole correction is partially small, it looks closer to a circle shape.

### v3

# Galaxy groups in the presence of Cosmological Constant: Increasing the Masses of Groups

@@ Abstract
The boundaries of galaxy groups and clusters are defined by the interplay between the Newtonian attractive force and the decoupling from the local expansion of the Universe. This work extends the definition of a zero radial acceleration surface (ZRAS) and the turnaround surface (TS) for a general distribution of the masses in an expanding background, governed by the cosmological constant. We apply these definitions to different galaxy groups in the local Universe, mapping these groups up to ten megaparsec distances. We discuss the dipole and the quadrupole rate for the Local Group of Galaxies and the implementations on the Hubble diagram correction and galaxy groups virialization. With these definitions, we present the surfaces showing the interplay between the local expansion vs the local Newtonian attraction for galaxy groups in the local Universe. Further, we estimate the masses of different galaxy groups and show that the inclusion of the Cosmological Constant in the analysis predicts these masses to be higher by 5-10%. For instance the Local Group of Galaxies is estimated to be $\left(2.47\pm 0.08\right)\cdot 10^{12}M_{\odot}$. For the groups with enough tracers close to the TS, the contribution of the Cosmological Constant makes the masses to be even higher. The results show the importance of including the local cosmic expansion in analyzing the Cosmic Flow of the local Universe.
@@ I Introduction
Dark energy is an unknown type of energy that influences the Universe on its largest scales Peebles and Ratra (2003). Its primary impact is to propel the accelerating expansion of the Universe. The first observational evidence for dark energy emerged from supernova measurements Scolnic et al. (2018) and later was confirmed by the Cosmic Microwave Background radiation measurements Aghanim et al. (2020) and the Baryon Acoustic Oscillations. Beyond its impact on large scales, dark energy exerts a significant influence on the Local Universe Courtois et al. (2013). Galaxy groups can be characterized by the radius of decoupling from cosmic expansion Tully (2015a). In an isolated system, the dark energy creates a local repulsion force resisting the Newtonian attraction:
@@
where $r$ is the separation, $M$ is the total mass of the binary system, $G$ is the Newtonian gravitational constant, $c$ is the speed of light and $\Lambda$ is the cosmological constant. It is possible to infer the cosmological constant from the cosmological parameters via: $\Lambda=3\Omega_{\Lambda}\left(H_{0}/c\right)^{2}\;$, where $H_{0}$ the Hubble parameter that estimates the current expansion rate, and $\Omega_{\Lambda}$ is the dark energy density rate.
Since the dark-energy contribution depends on the mass density, we define two useful surfaces: the Zero Radial Acceleration Surface (ZRAS) and the Turnaround Surface (TS). These two definitions are compatible for the spherical case, and in this paper we generalize them for a general density structure.
@@
    $\bar{r}_{0}\equiv\sqrt[3]{\frac{3GM}{\Lambda c^{2}}}\approx 1.11\,\text{Mpc}\,\left(\frac{M}{10^{12}\,M_{\odot}}\right)^{1/3}\left(\frac{67\,km/s/Mpc}{H_{0}}\right)^{2/3},$ (2)
    which is called the zero gravity radius. $M_{\odot}$ is the solar mass. There are different methods to measure the zero gravity radius through the local Hubble expansion Del Popolo and Chan (2022) or by estimating the points at which the sphere breaks away from the general expansion and reaches a maximum radius Pavlidou and Tomaras (2014); Faraoni et al. (2015); Faraoni (2016); Giusti and Faraoni (2021). For the spherical case, the zero gravity radius is sufficient, but for a general density distribution one has to use the ZRAS, which reads Dolgachev et al. (2003):
@@
  • Turnaround Surface (TS) or Zero Velocity Surface – In statistical mechanics, the virial theorem provides a general equation that relates the average over time of the total kinetic energy of a closed and stable system of discrete particles, bound by a conservative force, to the average of the potential energy of the system Lahav et al. (1991); Mota and van de Bruck (2004); Wang (2006); Chernin et al. (2012). The conventional virial theorem can be modified by an extra term due to the contribution of the dark energy, besides the Newtonian attraction. By imposing that all the kinetic energy of the system is positive ($K>0$) one gets the condition Nowakowski et al. (2002):
    $\frac{|U_{G}|}{2}<U_{\Lambda}.$ (4)
@@
    The condition for positive kinetic energy is extended for a general density case and gives the definition for the virialization surface. In the standard case where the dispersion of the velocities is known, the virial theorem gives the lowest possible radii for a closed system. In this case, it is possible to calculate the value of the average spherical radii. Consequently, there exists a surface that obeys the virial theorem for the particles inside the surface.
Galaxy groups can be characterized by the radius of decoupling from cosmic expansion  Tully (2015a). However, this definition has been made without the presence of dark energy. Ref. Tully (2015a, b); Kourkchi and Tully (2017) consider the real space and  Tempel et al. (2016, 2017); Sorce and Tempel (2018) consider the redshift space. In this letter we extend these definitions from a spherical symmetric case to a general density using multipole expansion and define the corresponding ZRAS and TS.
The structure of the paper is as follows: Section II reviews the theoretical framework for a spherical density model with dark energy. Section III extends the analysis for a galaxy pair. Section IV shows the applications for galaxies in the local Universe and maps the local galaxy complexes correspondingly. Section V generalizes the formulation for general density with multipole expansion. Section VI calculates the masses of certain groups of galaxies in the presence of the cosmological constant $\Lambda$. Section VII contains the conclusion and summarizes the results of the work.
@@ II Point Particle and Spherical Symmetry
Zero Acceleration Surface
@@
Turnaround Surface
The virial theorem provides a general equation that relates the average over time of the total kinetic energy of a stable system of discrete particles, bound by a conservative force, to that of the total potential energy of the system. The virial theorem relates the average of a potential energy with a power law of the form $U\sim r^{n}$, via the relation Landau and Lifschits (1975); Goldstein et al. (2001): $2\langle K\rangle=n\langle U\rangle$, where $K$ is the kinetic energy of the system, $U$ is the potential energy of the system, and the averages are taken over a long time. The average yields:
$2K+U_{G}=2U_{\Lambda},$ (7)
@@
$K=\frac{1}{2}\sum m_{i}v_{i}^{2},\quad U_{G}=G\sum_{i<j}m_{i}m_{j}/r_{ij},\quad U_{\Lambda}=\frac{\Lambda c^{2}}{6}\sum_{i}m_{i}r_{i}^{2}.$ (8)
@@
As a simple model for the galaxy groups in the local Universe, it is assumed that most of the mass is concentrated in the massive galaxies, and the other dwarf galaxies are test particles with mass $m$. For test particles obeying the Eqs. (8) with the condition (4), one finds $r<\bar{r}_{0}$, where $\bar{r}_{0}$ is the ZRAS for the massive galaxy $M$. It means that for a closed system within the zero radial acceleration surface. The estimate indicates that gravitation is stronger than repulsion within the volume of a group, making the existence of gravitationally bound systems with finite orbits possible. Based on this condition, it was suggested in Membrado and Pacheco (2016) to use surface term effects as mass estimators. From this inequality, it is clear that this gives an upper value for the mass inside the spherical surface.
III Galaxy Groups with Axial Symmetry
@@ III.1 The surfaces
where $m_{1,2}$ are the fractional masses from the total mass: $m_{1,2}=M_{1,2}/(M_{1}+M_{2})$. Fig. (1) shows an illustration of this system. A test particle around these galaxies feels the potential
@@
Figure 1: Stream lines for a galaxy pair with masses $M_{1,2}$ located at $\left(0,z_{1,2}\right)$. The stream lines shows the attractor for the gravitational force inside the surface described by the Eq. (16), vs the repulsion field outside. The zero acceleration surface is presented in red line.
Critical points
@@
$F_{z}=\frac{\Lambda c^{2}z}{3}-\frac{GM_{1}(z-z_{1})}{\left[\rho^{2}+(z-z_{1})^{2}\right]^{3/2}}-\frac{GM_{2}(z-z_{2})}{\left[\rho^{2}+(z-z_{2})^{2}\right]^{3/2}},$ (12a)
$F_{\rho}=\frac{\Lambda c^{2}\rho}{3}-\frac{GM_{1}\rho}{\left[\rho^{2}+(z-z_{1})^{2}\right]^{3/2}}-\frac{GM_{2}\rho}{\left[\rho^{2}+(z-z_{2})^{2}\right]^{3/2}}.$ (12b)
@@
  • Points A, C $(0,\pm z_{0})$: The force at these points is zero, ${\bf F}_{A,C}={0}$. For $\rho=0$, we get $F_{\rho}=0$ from Eq. (12b). Imposing $F_{z}=0$ in Eq. (12a) gives:
@@
    $\frac{GM_{1}}{\left(\rho_{0}^{2}+z_{1}^{2}\right)^{3/2}}+\frac{GM_{2}}{\left(\rho_{0}^{2}+z_{2}^{2}\right)^{3/2}}=\frac{\Lambda c^{2}}{3}.$ (14)
@@
Radial acceleration surface
@@
$\frac{\Lambda c^{2}r^{2}}{3}=GM_{1}\frac{r^{2}-zz_{1}}{\left(r^{2}+z_{1}(z_{1}-2z)\right)^{3/2}}+GM_{2}\frac{r^{2}-zz_{2}}{\left(r^{2}+z_{2}(z_{2}-2z)\right)^{3/2}}.$ (16)
Fig. (1) shows in dashed black line the solution for this curve. The plot shows that this curve differentiates the gravitational attraction from the global cosmic expansion.
Figure 2: Gradient lines (in blue) are for the Local Group of galaxies including the MW (with $M_{1}\approx 10^{12}\,M_{\odot}$) and M31 ($M_{2}\approx 2\cdot 10^{12}\,M_{\odot}$) with a distance of $0.77\,Mpc$, located at $(0,0.77\cdot 2/3)$ and $(0,-0.77/3)$, respectively. The center of mass is at the origin of the system of coordinates. The streamlines show the areas for the domination of the Newtonian attraction vs the cosmic repulsion. The dwarf galaxies around the MW and M31 are also added as points with the corresponding colour of the velocity from the LG center of mass. Left: The ZRAS is shown with dashed black lines vs the monopole case in red. Right: The TS is shown in purple vs the monopole case in red. Both definitions for group boundaries are close to the monopole case. The $\bar{r}_{0}$-surface for the monopole case is plotted as the dashed red circle in both panels.
Turnaround surfaces
@@ III.2 Expansion
To study the properties of the potential, one can expand the potential around $z_{i}/r\ll 1$, assuming that the observer is distant from the group. The approximation gives the following Peñarrubia et al. (2014):
@@
$\sigma_{\text{Quad}}$$\equiv$ $\frac{m_{1}}{2}z_{1}^{2}+\frac{m_{2}}{2}z_{2}^{2}.$
Since we work in CoM coordinates, i.e. CoM is at the origin (Eq. 9), the dipole of the system is zero Peñarrubia et al. (2016). The dipole, octupole, and all multipoles with odd powers in $r^{i}$ defined in the center of mass of any system should be equal to zero. In order to quantify the contribution of the quadrupole, we can test the dimensionless quantity
@@
IV Mapping the Close-by Groups
@@
MW Mass $[0.9,1.3]\cdot 10^{12}\,M_{\odot}$ Wang et al. (2020)
M31 Mass $[1.0,2.0]\cdot 10^{12}\,M_{\odot}$ Sawala et al. (2023a)
@@
Table (2) shows the data we consider for the LG of galaxies to estimate the characteristic surfaces. Table (2) shows the different values for the zero acceleration surface (monopole), the critical points ZRAS and TS. We can see from the figure that the surfaces is non-spherical, where one edge is larger than $\bar{r}_{0}$ and the other edge is smaller than $\bar{r}_{0}$. The intersection of this surface with the symmetry plane (as in Fig. (2)) is represented by a slightly asymmetric oval that lies between two circles, where $z_{1}/\bar{r}_{0}\approx 33\%$ and $z_{2}/\bar{r}_{0}\approx 20\%,$ are the partial corrections for the galaxy locations over the monopole ZRAS, with an error less than one percent. For the LG, the corresponding quadrupole correction is $\tilde{\sigma}_{\text{Quad}}^{(\text{LG})}<3.2\%$. Therefore, treating the quadrupole as a perturbation over the monopole and Dark Energy parts is a valid assumption.
@@
One Galaxy – The simplest case is one dominant giant galaxy and some dwarf galaxies around. The NGC 253 system and the IC-342 group and M101 galaxies seem to satisfies this condition. In this case, the spherical symmetry is a good approximation, unless there is a distribution of dark matter around the system that breaks this symmetry. Therefore, around these giant galaxies the TS looks as a perfect circle, despite that the dark matter distribution could give additional quadrupole corrections.
Galaxy Pair – The LG has two main components, the MW and M31. In fact, this binary structure has been suggested as a possible source for the ejection of dwarf galaxies. However, the zero-gravity surface calculated for the LG is almost spherically symmetric at the relevant distances. The CenA & M83 Complex system seems to be dominated by two giant galaxies: Centaurus A (NGC 5128) and M83 (NGC 5236), where the other galaxies are dwarf galaxies around these. Another pair is the M81-M82 Complex, where the M81 group is a galaxy group in the constellations Ursa Major and Camelopardalis. For these systems, the TS gives an egg shape as discussed in the LG, but since the quadrupole correction is partially small, it looks closer to a circle shape.
Figure 3: Two projections of the Local Group (up to 10 Mpc) and their immediate vicinity. The dashed curves show the gravitational bond (with dark energy) for massive galaxies. The figure includes the Local Group, the Cen-A/M83 complex, the M81-M82 group, and the NGC 253 Group. The colours of the dwarf galaxies around show the line of sight velocities as taken from Cosmic Flow Hoffman et al. (2023).